A teacher uses a geiger Muller tube and counter to measure the number of counts in 60 seconds for a radioactive rock.

the counter recorded 819 counts in 60 seconds. the background radiation count was 0.30 counts per second. Calculate the count rate for the rock
1
Expert's answer
2019-11-12T17:25:22-0500

The number of discharges due to the radioactive sample is 819 - 60*0.3 = 801, therefore the activity of the sample "A = \frac{801}{60} = 13.4 (Bq)"